# Web Privacy based on Contextual Integrity: Measuring the Collapse of Online Contexts

### v1

@@ Abstract.
The collapse of social contexts has been amplified by digital infrastructures but surprisingly received insufficient attention from Web privacy scholars. Users are persistently identified within and across distinct web contexts, in varying degrees, through and by different websites and trackers, losing the ability to maintain a fragmented identity. To systematically evaluate this structural privacy harm we operationalize the theory of Privacy as Contextual Integrity and measure persistent user identification within and between distinct Web contexts. We crawl the top-700 popular websites across the contexts of health, finance, news & media, LGBTQ, eCommerce, adult, and education websites, for 27 days, to learn how persistent browser identification via third-party cookies and JavaScript fingerprinting is diffused within and between web contexts. Past work measured Web tracking in bulk, highlighting the volume of trackers and tracking techniques. These measurements miss a crucial privacy implication of Web tracking - the collapse of online contexts. Our findings reveal how persistent browser identification varies between and within contexts, diffusing user IDs to different distances, contrasting known tracking distributions across websites, and conducted as a joint or separate effort via cookie IDs and JS fingerprinting. Our network analysis can inform the construction of browsers’ storage containers to protect users against real-time context collapse. This is a first modest step in measuring Web privacy as contextual integrity, opening new avenues for contextual web privacy research.
@@ 1. Introduction
Digital infrastructures have increased the scope and frequency of the collapse of social contexts (Frost-Arnold, 2021). It is now almost impossible for data subjects to maintain a fragmented identity online. This is especially evident on the Web, where ubiquitous tracking has been an inevitable part of the Web’s structure, fueling the 225 billion dollar online advertising industry (IAB, 2024). Surprisingly, even though the intentional collapse of multiple contexts into one has been so pervasive for Web users, it has not received sufficient scholarly attention. We lack a systematic measure of online context collapse on the Web.
Past empirical work studied web tracking in bulk, indirectly addressing context collapse. Scholars provided an essential window into the widespread uses of tracking practices but had not provided a nuanced comparative understanding of context collapse. Scholars highlighted the prevalence of certain trackers and the centralized nature of the advertising industry. Comparative findings on tracking across websites are sparse, mainly revealing invasive tracking by news websites along with lower amounts of tracking on public websites, based on different incentives to monetize content.
@@
To fill this gap, we operationalize the theory of Privacy as Contextual Integrity (CI). According to CI, profiling users based on information gleaned across social contexts, fails to respect contextual informational norms, and in so doing, violates people’ privacy expectations. Accordingly, we examine how persistent browser identification by stateful and stateless mechanisms is happening by third-party trackers, both within and between different online contexts. This enables to assess when and by whom the norms of the original context are being violated and context collapse is taking place.
We collected stateful and stateless browser identification patterns across top 700 popular websites for 27 days, in seven different online contexts – Health, Finance, Education, LGBTQ, Adult, News & Media, e-Commerce. We offer a close empirical examination, based on client-side interactions with websites, of the extent that trackers use persistent identifiers for users across and within social contexts. We assume, based on previous works (Olejnik et al., 2012; Bird et al., 2020a), that once a user ID is linked by trackers across different contexts, browsing history can be used to infer sensitive attributes about a person – health condition, desire to quit a job, political orientation, and etc. Even though we cannot observe how advertisers use data to decide on the best targeting method per user, we know that there is a natural incentive in this ecosystem to aggregate data in order to learn large fraction of user’s history and potential future behavior for targeting purposes (Bird et al., 2020b; Acar et al., 2014; McGuigan, 2019).
We show which online contexts are more vulnerable to single and multiple context collapse. We detail how the amount of tracking and participating websites varies between multiple and single context collapse, how the distribution of trackers across websites differs from the known, long tail distribution of trackers on the Web, and how far user IDs travel between online contexts. We also reveal the extent of overlap between the usage of cookie IDs and JS fingerprinting for context collapse. We then visualize the collapse of contexts via network graphs and calculate the associated chromatic numbers to inform the construction of browser’s containers and prevent real-time context collapse for users.
@@
  • Develops novel Web privacy measurement: We designed a robust method to identify persistent browser identification and the collapse of single and multiple contexts on the Web, offering a novel measurement of Web privacy through the operationalization of the theory of privacy as Contextual Integrity, which rarely gets attention from Web privacy scholars.
  • Reveals the structural features context collapse on the Web: We show, for the first time, how persistent browser identification varies within and between web contexts, highlighting the structural features that determine the level of real-time context collapse on the Web.
@@
In the next section we motivate our research by explaining our argument on how online contexts collapse on the Web and why related work has yet to systematically evaluate and assess this privacy violation. We then turn to operationalize Web privacy based on CI, discuss our measurement framework, data collection, analysis, limitations, results, contribution, and suggestions. We conclude with future research avenues and open questions on the contextual measurement of Web privacy.
@@ 2. The Collapse of Online Contexts on the Web
The collapse of online contexts has been mostly studied by social media and platform researchers, who demonstrated how these infrastructures blur public and the private, make content available beyond the temporal moment of its creation, and obscure the viewership and unlimited audience for online content (danah boyd, 2012; Vitak, 2012; Wesch, 2009). Surprisingly, context collapse on the Web and its associated privacy implications has not received sufficient scholarly attention and yet to be measured systematically.
@@
The intentional online collapse of multiple contexts into one audience regularly happens on the Web. As opposed to our offline interactions, where we barely take information out of its original context and would not imagine, for example, reading private messages over the shoulder of a stranger (Berjon, 2021), our online interactions, and specifically our interactions with websites, are full of intentional context collapse. Users’ Web history and site behavior seeps beyond the target website itself to external trackers for marketing purposes. Such context collision is accessible to marketers, often against users’ expectations (McClain et al., 2023), happening both within- and between- different web contexts, and enabling trackers to stitch together sensitive information about users’ queries and site interests from different web contexts. The majority of users are ‘not comfortable’ with such tracking and profiling, and do not want their data to be used for purposes other than proving the service they requested (McDonald and Cranor, 2010). They are especially uncomfortable with actors navigating their data between different contexts to get a better understanding of what users might be interested in (inmoment, 2018). Notice and consent mechanisms fail to properly protect users (Barocas and Nissenbaum, 2014; Martin, 2015), leaving them exposed to constant collapse of their distinct online contexts.
@@
Empirical Web privacy researchers have been studying trackers in bulk, inspecting interactions between users’ browsers and thousands of websites, with little sensitivity to the different contexts for users on the Web or to degrees of context collapse by the actors involved. Millions of third-party requests were examined by researchers to highlight the centralized nature of the advertising ecosystem and the significant presence of ‘top-trackers’ (e.g. Google, Meta, X) across the majority of measured traffic. The distribution of third-parties across the Web was found to have a long tail, with the majority of trackers operate on less than 1 percent of the Web. (Karaj et al., 2018; Yang and Yue, 2020; Englehardt and Narayanan, 2016; Solomos et al., 2020; Lerner et al., 2016; Libert, 2015; Libert and Binns, 2019; Sankin and Mattu, 2020; Bird et al., 2020a). Users’ web history collected by third-party trackers was found to be unique to users and useful for extracting insights about users (Olejnik et al., 2012). Researchers showed how users can be re-identified via their web history, with the dominant trackers - Google & Meta - unsurprisingly enjoying a privileged position in the ability to do so (Bird et al., 2020a).
@@
Scholars who did consider different categories of websites in their results show variance in the amount of tracking within categories. An overarching finding from these efforts is that news websites enable more third-party tracking than other types of websites, as opposed to public websites that demonstrate less presence of trackers (Englehardt and Narayanan, 2016; Libert and Binns, 2019; Urban et al., 2020; Karaj et al., 2018; Yang and Yue, 2020; Lerner et al., 2016). The difference in the amount of tracking for each category is usually associated with the incentives for publishers to include third-party trackers. Different business models and funding resources across websites are hypothesized as a possible explanation for the observed variance of tracking amount across website categories (Englehardt and Narayanan, 2016).
@@
Empirically, we diverge from existing research by pivoting our observations of website interconnectedness based on distinct contexts and spot patterns of persistent identification that go beyond common presence of a third-party among first-party websites, since not all third-party trackers are participating in persistent user identification.
Theoretically, past works leave us with many questions on context collapse on the Web. Existing studies do not distinguish persistent browser identification patterns within and between Web contexts and the extent to which they enable the collapse of online contexts. Which third-parties contribute to the collapse of online contexts? Who is responsible for context collapse beyond the ’usual suspects’ of advertising giants? What are the websites that are more or less vulnerable to online context collapse? We lack a comparative understanding across popular web contexts of how using persistent user identifiers differs among contexts. For example, do users’ interactions with popular healthcare websites are more/less commonly linked to users’ browsing behavior in other popular web contexts? To what extent the fact that users browse to consume adult content can serve advertisers when the same users engage in e-commerce, seek financial advice, or look for health support? We aim to conduct a context-sensitive analysis of persistent browser identification across the Web and gain a better understanding of how and by whom trackers collapse users’ online contexts.
Alarmingly, context collapse has become a pervasive feature of our online lives, shaping what is known about us in the world and forms our epistemic practices, habits, and opportunities (Frost-Arnold, 2021). The tendency to collapse online contexts risks a main feature enabled by privacy - the ability of individuals to hold multiple identities in different social contexts. Identity is by nature fragmented: we present ourselves as different people in different contexts (Berjon, 2021). The collapse of online contexts, however, leads to overlapping individual identities through the intermingling of distinct networks.
To measure the collapse of online contexts on the Web, and systematically understand how the third-party structure of the Web has been enabling these privacy harms, we turn to the theory of privacy as CI and develop a novel measurement of Web privacy.
@@ 2.1. Web Privacy as Contextual Integrity
Based on the category of websites, we can assess their context and evaluate the purpose of user engagement, tagging users’ web history and site interactions to a context. It is a violation of privacy to share information outside original web contexts in ways that do not advance the purpose for which these contexts exist for users. Notice and consent choices are unable to capture the ways users’ web history and site interaction data lead to the collapse of online contexts because actual data flows on the Web are beyond individual cognitive capacity (Barocas and Nissenbaum, 2014; Martin, 2015).
In times, violating contextual norms is virtuous since it advances a greater socially desired goal. Cases of whistleblowing, or the sharing of health information to mitigate a global pandemic can all be desired collapses of online contexts. But the collapse of contexts for user targeting and marketing purposes goes against users’ expectations and does not respect the norms in the original Web context users chose to engage in. According to past surveys, users and consumers expect their online information flows to maintain the integrity of the context in which they operate (McClain et al., 2023).
Contextual Integrity flags violation of contextual norms in two distinct browsing scenarios: (1) within a web context, when, for example users seek health advice across an array healthcare websites, persistent user identification by third-party trackers enables the construction of rich user profiles based on users’ Web history across sites from the same context. These users’ site histories are revealing a great deal of users’ health concerns, and collecting them by an external third-party tracker for marketing purposes is collapsing the online context for users interacting with this array of healthcare sites. We refer to this as ’single context collapse. (2) Between contexts, persistent browser identification further changes the meaning of sensitive information. It enables trackers to bring together users’ web history and interests from different online contexts, collapsing not just one, but two or more online contexts for the sake of marketing and advertising. For example, when users’ browsers are persistently identified during subsequent visits to healthcare and LGBTQ websites, users experience context collapse in both, allowing trackers and marketers to build user profiles based on user engagement in both contexts, without advancing the original purpose of these contexts. We refer to this as ’multiple contexts collapse.’
@@
We use CI’s contextual factors to map and label the information flows we wish to assess, as detailed in figure 1. To measure multiple contexts collapse we inspect information flows in which: (1) the sender of the information is the user’s browser; (2) the recipient of the information is tracker A in context Y; (3) the information type is the assigned user ID by tracker A in context X; (4) the transmission principle is notice & consent; and (5) the data subject is the individual who interacts with the website through the browser.
To measure within-context, or single context collapse we inspect information flows in which: (1) the sender of the information is the user’s browser; (2) the recipient of the information is tracker A in Website 2 within context X; (3) the information type is the assigned user ID by tracker A in Website 1 within context X; (4) the transmission principle is notice & consent; and (5) the data subject is the individual who interacts with the website through the browser.
Figure 1. Information flows that capture multiple (between) context collapse and single (within) context collapse
@@ 3.1. Persistent User Identification on the Web
We pivot our empirical observations around both third-party identification cookies and JavaScript fingerprinting functions that assign a dynamic ‘advertising identifier’ to users as they browse the Web. Those ‘advertising Identifiers’ are a key prerequisite for the industry to address a user for personalized/targeted advertising since all user-centric data is associated with it (Europe, 2020). Despite the default block of third-party cookies by several browsers, Google, that operates the browser with the largest market share (Statista, 2024), has recently announced that third-party cookies are here to stay, making the collapse on online contexts via cookie IDs a viable object of study (Spadafora, 2024).

### v2

@@ Abstract.
The collapse of social contexts has been amplified by digital infrastructures but surprisingly received insufficient attention from Web privacy scholars. Users are persistently identified within and across distinct Web contexts, in varying degrees, through and by different websites and trackers, losing the ability to maintain a fragmented identity. To systematically evaluate this structural privacy harm, we operationalize the theory of Privacy as Contextual Integrity and measure persistent user identification within and between distinct Web contexts. We crawl the top-700 popular websites across the contexts of health, finance, news & media, LGBTQ, eCommerce, adult, and education websites, for 27 days, and created network graphs to learn how persistent browser identification via third-party cookies and JavaScript fingerprinting is diffused within and between Web contexts. Past work measured Web tracking in bulk, highlighting the volume of trackers and tracking techniques. These measurements miss a crucial privacy implication of Web tracking - the collapse of online contexts. Our findings reveal how persistent browser identification varies between and within contexts, diffusing user IDs to different distances, contrasting known tracking distributions across websites, and conducted as a joint or separate effort via cookie IDs and JS fingerprinting. Our network analysis informs the construction of browsers’ storage containers to protect users against real-time context collapse. This is a first modest step in measuring Web privacy as Contextual Integrity, opening new avenues for contextual Web privacy research.
@@ 1. Introduction
Digital infrastructures have increased the scope and frequency of the collapse of social contexts (Frost-Arnold, 2021). It is now almost impossible for data subjects to maintain a fragmented identity online. This is especially evident on the Web, where ubiquitous tracking has been an inevitable part of the Web’s structure, fueling the 225 billion dollar online advertising industry (IAB, 2024). Surprisingly, even though the intentional collapse of multiple contexts into one has been so pervasive for Web users, it has not received sufficient scholarly attention. We lack a systematic measure of Web context collapse.
Past empirical work studied Web tracking in bulk, indirectly addressing context collapse. Scholars provided an essential window into the widespread uses of tracking practices, but had not provided a nuanced comparative understanding of context collapse. Scholars highlighted the prevalence of certain trackers and the centralized nature of the advertising industry. Comparative findings on tracking across websites are sparse, mainly revealing invasive tracking by news websites along with lower amounts of tracking on public websites, based on different incentives to monetize content.
@@
To fill this gap, we operationalize the theory of Privacy as Contextual Integrity (CI). According to CI, profiling users based on information gleaned across social contexts, fails to respect contextual informational norms, and in so doing, violates people’ privacy expectations. Accordingly, we examine how persistent browser identification by stateful and stateless mechanisms is happening by third-party trackers, both within and between different online contexts. We assess when and by whom the norms of the original context are being violated and context collapse is taking place.
We collected stateful and stateless browser identification patterns across top 700 popular websites for 27 days, in seven different online contexts – Health, Finance, Education, LGBTQ, Adult, News & Media, e-Commerce. We offer a close empirical examination, based on client-side interactions with websites, of the extent that trackers use persistent identifiers for users across and within contexts. We assume, based on previous works (Olejnik et al., 2012; Bird et al., 2020a), that once a user ID is linked by trackers across different contexts, browsing history can be used to infer sensitive attributes about a person – health condition, desire to quit a job, political orientation, and etc. Even though we cannot observe how advertisers use data to decide on the best targeting method per user, we know that there is a natural incentive in this ecosystem to aggregate data in order to learn large fraction of user’s history and potential future behavior for targeting purposes (Bird et al., 2020b; Acar et al., 2014; McGuigan, 2019).
We show which online contexts are more vulnerable to single and multiple context(s) collapse. We detail how the amount of trackers and participating websites varies between multiple and single context collapse, how the distribution of trackers across websites differs from the known, long tail distribution of trackers on the Web, and how far user IDs travel between online contexts. We also reveal the extent of overlap between the usage of cookie IDs and JS fingerprinting for context collapse. We then visualize the collapse of contexts via network graphs and calculate the associated chromatic numbers to inform the construction of browser’s cookie storage containers and prevent real-time context collapse for users.
@@
  • Develops a novel Web privacy measurement: We designed a robust method to identify persistent browser identification and the collapse of single and multiple contexts on the Web, offering a novel measurement of Web privacy through the operationalization of the theory of privacy as Contextual Integrity, which rarely gets attention from Web privacy scholars.
  • Reveals the structural features of context collapse on the Web: We show, for the first time, how persistent browser identification varies within and between Web contexts, highlighting the structural features that determine the level of real-time context collapse on the Web.
@@
In the next section we motivate our research by explaining our argument on how online context collapse takes place on the Web and why related work has yet to systematically evaluate and assess this privacy violation. We then turn to operationalize Web privacy based on CI, discuss our measurement framework, data collection, analysis, limitations, results, contribution, and suggestions. We conclude with future research avenues and open questions on contextual measurement of Web privacy.
@@ 2. The Collapse of Online Contexts on the Web
The collapse of online contexts has been mostly studied by social media and platform researchers, who demonstrated how these infrastructures blur the public and the private, make content available beyond the temporal moment of its creation, and obscure the viewership and unlimited audience for online content (danah boyd, 2012; Vitak, 2012; Wesch, 2009). Surprisingly, context collapse on the Web and its associated privacy implications has not received sufficient scholarly attention and yet to be measured systematically.
@@
The intentional collapse of multiple contexts into one audience regularly occurs on the Web. Unlike our offline interactions, where we barely take information out of its original context and would not imagine, for example, reading private messages over the shoulder of a stranger (Berjon, 2021), our online interactions, and specifically our interactions with websites, are full of intentional context collapse. Users’ Web history and site behavior seeps beyond the target website itself to external trackers for marketing purposes. Such context collision is accessible to marketers, often against users’ expectations (McClain et al., 2023), happening both within- and between- different Web contexts, and enabling trackers to stitch together sensitive information about users’ queries and site interests from different web contexts. The majority of users are ‘not comfortable’ with such tracking and profiling, and do not want their data to be used for purposes other than proving the service they requested (McDonald and Cranor, 2010). They are especially uncomfortable with actors navigating their data between different contexts to get a better understanding of what users might be interested in (inmoment, 2018). Notice and consent mechanisms fail to properly protect users (Barocas and Nissenbaum, 2014; Martin, 2015), leaving them exposed to constant collapse of their distinct online contexts.
@@
Empirical Web privacy researchers have been studying trackers in bulk, inspecting interactions between users’ browsers and thousands of websites, with little sensitivity to the different contexts for users on the Web or to degrees of context collapse by the actors involved. Millions of third-party requests were examined by researchers to highlight the centralized nature of the advertising ecosystem and the significant presence of ‘top-trackers’ (e.g. Google, Meta, X) across the majority of measured traffic. The distribution of third-parties across the Web was found to have a long tail, with the majority of trackers operate on less than 1 percent of the Web (Karaj et al., 2018; Yang and Yue, 2020; Englehardt and Narayanan, 2016; Solomos et al., 2020; Lerner et al., 2016; Libert, 2015; Libert and Binns, 2019; Sankin and Mattu, 2020; Bird et al., 2020a). Users’ Web history collected by third-party trackers was found to be unique to users and useful for extracting insights about users (Olejnik et al., 2012). Researchers showed how users can be re-identified via their Web history, with the dominant trackers - Google & Meta - unsurprisingly enjoying a privileged position to do so (Bird et al., 2020a).
@@
Scholars who did consider different categories of websites showed variance in the amount of tracking within categories. An overarching finding was that news websites enable more third-party tracking than other types of websites, as opposed to public websites that demonstrate less presence of trackers (Englehardt and Narayanan, 2016; Libert and Binns, 2019; Urban et al., 2020; Karaj et al., 2018; Yang and Yue, 2020; Lerner et al., 2016). The difference in the amount of tracking for each category is usually associated with the incentives for publishers to include third-party trackers. Different business models and funding resources across websites are hypothesized as a possible explanation for the observed variance of tracking amount across website categories (Englehardt and Narayanan, 2016).
@@
Empirically, we diverge from existing research by pivoting our observations of website interconnectedness based on distinct contexts, and spot patterns of persistent identification that go beyond common presence of a third-party among first-party websites, since not all third-party trackers are participating in persistent user identification.
Existing studies do not distinguish persistent browser identification patterns within and between Web contexts and the extent to which they enable the collapse of online contexts. In contrast, our approach enables to engage with questions such as Which third-parties contribute to the collapse of online contexts? Who is responsible for context collapse beyond the ’usual suspects’ of advertising giants? What are the websites that are more or less vulnerable to online context collapse? We lack a comparative understanding across popular web contexts of how using persistent user identifiers differs among contexts. For example, do users’ interactions with popular healthcare websites are more/less commonly linked to users’ browsing behavior in other popular Web contexts? To what extent the fact that users browse to consume adult content can serve advertisers when the same users engage in e-commerce, seek financial advice, or look for health advice? We aim to conduct a context-sensitive analysis of persistent browser identification across the Web and gain a better understanding of how and by whom users’ contexts are collapsing.
Alarmingly, context collapse has become a pervasive feature of our online lives, shaping what is known about us and forms our epistemic practices, habits, and opportunities (Frost-Arnold, 2021). The tendency to collapse online contexts risks a main feature enabled by privacy - the ability of individuals to hold multiple identities in different social contexts. Identity is by nature fragmented: we present ourselves as different people in different contexts (Berjon, 2021). The collapse of online contexts, however, leads to overlapping individual identities through the intermingling of distinct networks.
To measure the collapse of online contexts on the Web, and systematically understand how the third-party structure of the Web has been enabling these privacy harms, we turn to the theory of privacy as Contextual Integrity (CI) and develop a novel measurement of Web privacy.
@@ 2.1. Web Privacy as Contextual Integrity
Based on the category of websites, we can assess their context and evaluate the purpose of user engagement, tagging users’ web history and site interactions to a context. It is a violation of privacy to share information outside original web contexts in ways that do not advance the purpose for which these contexts exist for users. Notice and consent choices are unable to capture the ways users’ Web history and site interaction data lead to the collapse of online contexts because the complex data flows for content monetization on the Web are beyond individual cognitive capacity (Barocas and Nissenbaum, 2014; Martin, 2015).
In times, violating contextual norms is virtuous since it advances a greater socially desired goal. Cases of whistleblowing, or the sharing of health information to mitigate a global pandemic can all be desired collapses of online contexts. But the collapse of contexts for user targeting and marketing purposes goes against users’ expectations and does not respect the norms in the original Web context. According to past surveys, users and consumers expect their online information flows to maintain the integrity of the context in which they operate (McClain et al., 2023).
Contextual Integrity is used to flag violation of contextual norms in two distinct browsing scenarios: (1) within a web context, when, for example users seek health advice across an array of healthcare websites, and persistent user identification by third-party trackers enables the construction of rich user profiles based on users’ Web history across sites from the same context. These users’ site histories are revealing a great deal of users’ health concerns, and collecting them by an external third-party tracker for marketing purposes is collapsing the online context for users interacting with the array of healthcare sites. We refer to this as ’single context collapse.’ (2) Between contexts, persistent browser identification further changes the meaning of sensitive information. It enables trackers to bring together users’ web history and interests from different online contexts, collapsing not just one, but two or more online contexts for the sake of marketing and advertising. For example, when users’ browsers are persistently identified during subsequent visits to healthcare and LGBTQ websites, users experience context collapse in both, allowing trackers and marketers to build user profiles based on user engagement in both contexts, without advancing the original purpose of these contexts. We refer to this as ’multiple contexts collapse.’
@@
We use CI’s contextual factors to map and label the information flows we wish to assess, as detailed in Figure 1. In that figure, the interaction of the user with the first context is on the left, followed by the interaction of the user with another context on the right. Specifically, to measure multiple contexts collapse we inspect information flows in which: (1) the sender of the information is the user’s browser; (2) the recipient of the information is tracker A in context Y; (3) the information type is the assigned user ID by tracker A in context X; (4) the transmission principle is notice & consent; and (5) the data subject is the individual who interacts with the website through the browser.
To measure within-context collapse we inspect information flows in which: (1) the sender of the information is the user’s browser; (2) the recipient of the information is tracker A in Website 2 within context X; (3) the information type is the assigned user ID by tracker A in Website 1 within context X; (4) the transmission principle is notice & consent; and (5) the data subject is the individual who interacts with the website through the browser.
Figure 1. Information flows that capture multiple and single context(s) collapse
@@ 3.1. Persistent User Identification on the Web
We pivot our empirical observations around both third-party identification cookies and JavaScript fingerprinting functions that assign a dynamic ‘advertising identifier’ to users as they browse the Web. Those ‘advertising Identifiers’ are a key prerequisite for the industry to address a user for personalized/targeted advertising since all user-centric data is associated with it (Europe, 2020). Despite the default block of third-party cookies by several browsers, Google, that operates the browser with the largest market share (Statista, 2024), has recently announced that third-party cookies are here to stay, making the collapse of online contexts via cookie IDs still very relevant (Spadafora, 2024).